### v2

# Towards the Terminator Economy: Assessing Job Exposure to AI through LLMs

Abstract
AI and related technologies are reshaping jobs and tasks, either by automating or augmenting human skills in the workplace. Many researchers have been working on estimating if and to what extent jobs and tasks are exposed to the risk of being automatized by AI-related technologies. Our work tackles this issue through a data-driven approach by: (i) developing a reproducible framework that uses cutting-edge open-source large language models to assess the current capabilities of AI and robotics in performing job-related tasks; (ii) formalizing and computing a measure of AI exposure by occupation, the TEAI (Task Exposure to AI) index, and a measure of Task Replacement by AI, the TRAI index, both validated through a human user evaluation and compared with the state of the art. Our results show that the TEAI index is positively correlated with cognitive, problem-solving and management skills, while it is negatively correlated with social skills. Applying the index to the US, we obtain that about one-third of US employment is highly exposed to AI, primarily in high-skill jobs, requiring a graduate or postgraduate level of education. We also find that AI exposure is positively associated with both employment and wage growth in 2003-2023, suggesting that AI has an overall positive effect on productivity. Considering specifically the TRAI index, we find that even in high-skill occupations, AI exhibits high variability in task substitution, suggesting that AI and humans complement each other within the same occupation, while the allocation of tasks within occupations is likely to change.
All results, models, and code are freely available online to allow the community to reproduce our results, compare outcomes, using our work as a benchmark to monitor AI’s progress over time.
Keywords:
AI, Large Language models, employment, skills.
1 Introduction
The famous 1984 movie “The Terminator” is set in a dystopian future where intelligent machines, created by a military defense system known as Skynet, become self-aware and perceive humanity as a threat, initiating a war to eliminate humans. Skynet creates advanced humanoid robots called Terminators to hunt down and kill human survivors. The Terminator has advanced learning algorithms that enable it to adapt to any environment, making it a formidable antagonist for humans. The debate and concerns about the impact of AI are often framed against the backdrop of the film’s setting. This paper takes these concerns seriously by developing an AI-centered assessment of the potential exposure of different occupations to artificial intelligence. Assessing the impact of technology on the labor market is not easy, as there are several potential channels at work. As stressed by (Acemoglu and Restrepo, 2019), technology has three major effects on labor demand. The first is the productivity effect, which operates through lower production costs brought about by new technologies. The second is the displacement effect of workers by machines and the like. These two effects operate in different directions and depend on whether technology substitutes for or complements human labor; in economic jargon, they depend on the elasticity of substitution between tasks. In addition, there is a third effect of technology: the creation of new tasks and activities in which labor can be productively employed (the reinstatement effect). Indeed, if we look at history, the reinstatement effect has been a central feature of all technological revolutions, which have continually created new opportunities for labor. For the reinstatement effect to take place, technology must have an impact wider than its narrow scope, with spillover effects in sectors/areas other than those for which it was designed. In other words, technology must have the features of a general purpose technology, which, according to (Lipsey et al., 2005), are pervasiveness across the economy, ability to generate complementary innovations, and improvement over time. AI provides strong arguments for being considered a general-purpose technology due to its broad applicability, potential productivity gains, and possibility to drive further innovation. However, these characteristics of AI create a relevant measurement problem, as it is extremely difficult to identify all the channels through which it affects the economy. This paper contributes to this field by developing a methodology for assessing AI exposure using Large Language Models (LLMs), with a granular approach that analyzes exposure for each task within each occupation. Our approach allows us to improve our understanding of the effect of AI on the labor market in several areas. Methodologically, we design and implement a reproducible framework to assess to what extent existing AI and robotics technologies can perform job-related tasks, relying LLMs. Rather than assessing AI exposure through external benchmarks such as expert judgment or AI patents and innovations data, we construct an internal assessment using LLM’s own evaluation. In other words, LLMs assess their ability to perform each of the 19K job-related tasks associated with about 1k occupations as described by O*NET [^1] O*NET: Occupational Information Network, is a comprehensive database containing detailed information on hundreds of standardized and occupation-specific descriptors. It is sponsored by the U.S. Department of Labor/Employment and Training Administration. O*NET serves as a resource to provide information on skills, abilities, knowledge, work activities, and interests associated with occupations. This approach has two main advantages. On the one hand, it is very granular. Each individual task, in fact, receives a specific assessment and a motivation by each LLM. On the other hand, it is fully transparent. Not only does it allow full reproducibility, [^2] All codes are available on GitHub at https://github.com/Crisp-Unimib/Terminator-Economy but more importantly, it allows comparability over time. AI is an innovative and fast-growing field where new models are expected to significantly outperform existing ones. New LLMs can be applied to our approach to measure the extent of their improvement in task performance. Consider taxi drivers, often used as an example of a potentially exposed occupation. The current assessment of the task of driving vehicles by LLMs is of average exposure, recognizing that current technologies can replace some tasks but not others that need to be performed by humans. However, with advances in AI systems, it is likely that the same activity will be fully automated in a few years’ time. New models will therefore assign a higher score and allow the extent of the change in exposure to be measured. From an economic perspective, we develop an AI exposure measure for each O*NET occupation based on specific assessment of individual tasks . Our AI exposure index is positively correlated with cognitive, problem-solving and management skills, consistent with the evidence that AI affects management and decision-making tasks; on the other hand, AI exposure is negatively correlated with social skills, where AI has clear limitations. We also develop a version of the index which is specifically constructed for measuring task substitutability by AI and which can be used to analyze the relationship between AI exposure and AI task substitutability. This analysis is enhanced by the fact that our approach provides, for each task, a description of AI involvement which constitutes a valuable wealth of information that can be used to analyze the human-machine interaction. Bringing our measure to the data, we show that, in the US, approximately 1/3 of employment is highly exposed to AI technologies; most of this employment (88%) is in high-skill jobs, requiring graduate or postgraduate level of education. In terms of labor market outcomes, using 4-year rolling regressions, we find that AI exposure is positively associated with both employment and wage growth in the period 2003-2023, suggesting that AI has an overall positive effect on productivity. The remainder of the paper is structured as follows: section 2 discusses the related literature, section 3 describes the methodology and the construction of the AI index, section 5 presents the results; finally, section 6 concludes.
2 Background and Related Works
AI and jobs
Since the seminal paper by (Autor et al., 2003), the task approach has proved very effective in analyzing the impact of technology on jobs. It divides work activities into tasks, each of which can be performed by humans or by machines. In this way, the distinction between capital and labor tasks is more precise, flexible, and able to shift over time. Indeed, capital and machines can substitute for labor in the performance of some tasks, while complementing it in the performance of others.
The task approach has been applied to analyze the effect of technology and trade (offshoring) (Acemoglu and Autor, 2011), to the long run effect of technology (Consoli et al., 2023) and to skill-task interaction (Colombo et al., 2019).
More recently this approach has been used to measure occupational exposure to computers and robots. In a seminal paper, (Frey and Osborne, 2017) estimated that up to 47% of jobs in the US are at risk of automation. [^3] See also (Nedelkoska and Quintini, 2018) for a similar approach. Subsequently, other attempts focused on developing measures of exposure to machine learning and robotics (Brynjolfsson and Mitchell, 2017; Acemoglu and Restrepo, 2020) and to AI (Felten et al., 2021; Webb, 2023; Eloundou et al., 2024; Pizzinelli et al., 2023).
Overall, these works find an extensive share of employment exposed to AI; the specific impact on occupations varies depending on the nuances that different indicators of the impact of technology capture, i.e., whether they focus on aspects of technology that impact more routine-based activities (Frey and Osborne, 2017) or more cognitive elements (Felten et al., 2021). A common feature of these papers is the attempt to quantify AI exposure through an external benchmark, which may be expert judgment or data analysis on patents and innovations. In contrast, our approach is based on an internal assessment, where LLM systems are asked to assess the suitability of tasks for AI. This approach has two main advantages. First, it is fully transparent, with results and outcomes being fully disclosed. Second, the approach is fully reproducible. This means that whenever new generations of LLM are available, they can be used in our approach to measure the change in task exposure that they imply.
Large Language Models
LLMs are powerful computational models designed to understand and generate human-like text by harnessing vast amounts of textual data, and have taken Natural Language Processing (NLP) by storm, achieving state-of-the-art performance in many tasks (Min et al., 2023). Typically, these models are based on Transformer architecture (Vaswani et al., 2017), powered by Attention mechanism (Luong et al., 2015; Bahdanau et al., 2014) and are composed by decoder-only stack.
These models are initially trained on an Autoregressive task Radford et al. (2018), where given a sequence of words $S=(w_{1},w_{2},...,w_{n-1})$ the training objective is to maximize the log-likelihood
$\sum_{i}\log P(w_{i}|w_{1},w_{2},....,w_{i-1};\theta^{T})$ where $\theta^{T}$ are the model parameters, in order to predict the next word in the sequence $\prod_{i=1}^{n}P(w_{i}\mid w_{1},...,w_{i-1})$. After being pre-trained, these models are fine-tuned for several tasks, providing examples of Natural Language Inference (Radford et al., 2018). Thanks to their capability to learn from context, known as in-context learning (Radford et al., 2019), LLMs can accomplish specific tasks with high accuracy (Zhao et al., 2023), exploiting prompt engineering methodologies such as zero-shot (Wei et al., 2021) and few-shot learning (Brown et al., 2020). In February 2025, Anthropic published a study [^4] https://www.anthropic.com/news/the-anthropic-economic-index to assess AI’s impact on jobs and tasks using O*NET, mapping millions of Claude conversations to O*NET tasks. By contrast, our approach (i) maps every O*NET occupation and task, (ii) employs a panel of three open LLMs to ensure reproducibility and minimize hallucinations, and (iii) allows the community to estimate the AI impact on other third-party taxonomies.
3 Building the AI Exposure Index
Our method can be summarized as follows: First, we obtain from O*NET the description of each task associated with each SOC occupation. Second, we apply LLMs to the task descriptions to obtain a rating of how well AI technologies can perform each task. Third, we aggregate the rating at the occupation level to obtain an AI exposure score. Finally, we apply our score to US data to assess the extent of AI exposure in the US labor market and the impact of AI on employment and wages. Figure 1 provides a graphical representation of our approach.
Figure 1: Graphical overview of the framework to compute the TEAI Index
3.1 Step 1: Compute the AI rate
In order to limit the risk of being driven by the well-known problem of LLMs called "hallucinations" (Ji et al., 2023), we design a framework involving three different LLMs aimed at identifying and limiting the false information generated, creating a consensus system between them.
Model choice
To ensure the reproducibility of this work we use three of the best-performing open source models, according to performance benchmarks, available on the open LLM rankings. [^5] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard The three selected models are Mistral 7B Instruct v 0.2 (Jiang et al., 2023), openchat 3.5 0106 (Wang et al., 2023) and orca mini v3 7b (Mukherjee et al., 2023).
Prompt design
The starting point is the ONET taxonomy which identifies 19281 tasks for 923 SOC occupations. [^6] We use the ONET 28.2 version released in February 2024. We formulate a five-shot prompt using the few-shot learning approach (Brown et al., 2020). We use each individual task description assigned to an occupation to ask the models how well, on a scale of 1 to 5, [^7] How well an AI system, which can be an LLM, Image Processing System or a Robot, could perform in the task on a scale of 1 to 5 where 1 stands for poor and 5 stands for excellent? the combination of different AI technologies could perform the input task and provide a discursive motivation for the evaluation. As AI technologies, we consider i) LLMs for textual data understanding, ii) Image Processing Systems for elaboration and decision-making based on visual data analysis, and iii) Robotic systems for physical execution. At the end of the prompt, the model is provided with five examples of the task to obtain more contextual and accurate results.
Model results
We iterate this process for each task provided by ONET and for each model chosen, ending up with three scores and natural language motivations provided by each model. Table 1 provides an example of the results after this stage for a selection of occupations and tasks.
O*NET Task Mistral Orca-mini Open-Chat Motivation provided by Mistral Direct or coordinate the supportive services department of a business, agency, or organization. 4 4 3 Robotics can automate repetitive tasks such as data entry, document scanning, and mail sorting. Image Processing Systems can analyze documents and emails for information extraction and categorization. LLMs can handle complex tasks like scheduling, coordinating resources, and communicating with team members. Combining these AI technologies can significantly improve the efficiency and accuracy of administrative services, allowing managers to focus on more strategic tasks. Conduct classes to teach procedures to staff. 2 3 2 LLMs could potentially create instructional materials, but they lack the ability to engage and assess staff performance during classes. Image Processing Systems have no role in this task. Robotics could theoretically conduct classes through pre-recorded videos, but the human interaction and customization required for effective training make this a poor solution.
Table 1: Extraction of two tasks for Administrative Services Managers occupation O*NET, enriched as discussed in Step1 to derive to what extent existing AI and robotics technologies can perform job-related tasks according to Mistral, Orca-mini, and Open-Chat LLMs. The rating scale expressed by the LLMs corresponds to the following categorical labels: 1: Poor, 2: Fair, 3: Average, 4: Good, 5: Excellent. Key activities are highlighted for a better reading
Evaluating consensus among LLMs
As mentioned above, the decision to use three different models was made to limit hallucinations. In order to construct a single indicator, we took a conservative approach by assigning to each task the value of the rating with the highest frequency among the three models; if the three rates were different, we chose the lowest.
To assess the agreement between the three rates expressed by the LLMs, for each single task we compute a consensus metric (Tastle and Wierman, 2007).
$Cns(a_{i})=1+\sum_{k=1}^{m}p_{k}\log_{2}(1-\frac{\lvert LV_{k}-\mu_{LV}\rvert}{d_{LV}})$ (1)
Equation (1) shows the consensus calculation in which, given an input task $a_{i}$, for which three rates are expressed by three different LLMs, $LV_{k}$ denotes the rate (on a scale from 1 to 5) expressed by the $k$-th LLM, $p_{k}$ denotes the relative frequency associated with $LV_{k}$; $\mu_{LV}$ defines the weighted average of the three $LV$ rating values computed using $p_{k}$ probabilities as weights; finally $d_{LV}$ represents the range of the rating scale adopted.
Specifically, for each $k$-th rate from 1 to 3 (corresponding to the three available rates), the absolute difference between the rating value $LV_{k}$ and the weighted average $\mu_{LV}$ is determined. This difference is then normalized by dividing by the scale size $d_{LV}$. The logarithmic function $\log_{2}$ is applied in order to moderate the impact of larger discrepancies from the mean.
This process allows the three rates to be combined into a single measure of agreement. A higher consensus value indicates greater agreement between the observed rating values expressed for each task by the three different LLMs.
Analogously, to estimate the similarity between the motivations provided by the LLMs, we compute the centroid of semantic cosine similarity (Rahutomo et al., 2012), between them. The embedding vectors for the centroid computation are obtained using an open source Transformer model: as for the LLMs, we chose the Transformer model to be used in accordance with the Massive Text Embedding Benchmark (MTEB) Leaderboard by (Muennighoff et al., 2023). [^8] https://huggingface.co/spaces/mteb/leaderboard. Having English-language motivations, the choice fell on the UAE-Large-V1 https://huggingface.co/WhereIsAI/UAE-Large-V1, which represented an excellent compromise between effectiveness and efficiency, given its small size.
Table 2 shows the results following the calculation of semantic similarity and consensus between the different rates. In particular, since higher cosine similarity values reflect higher semantic similarities between the text of the LLM motivations, we expect a strong correlation between the consensus metric and cosine similarity. The heat map represented in figure 2 shows that both the values of cosine similarity and the consensus metric are extremely high, with an average close to 0.9 in both cases.
On the one hand, this suggests coherence between LLM-generated rates and the associated motivations, and on the other hand, it adds robustness to our conservative approach in choosing between different models.
The result of this process is a single score te that returns a metric from 1 to 5 measuring the extent to which AI can perform each specific task, and a quantitative indicator of the similarity between the discursive motivations generated by the models.
O*NET Occupation O*NET Task Similarity Consensus
Administrative Services Managers Direct or coordinate the supportive services department of a business, agency, or organization. 0.918 0.828
Conduct classes to teach procedures to staff. 0.935 0.828
Taxi Drivers Complete accident reports when necessary. 0.948 0.828
Vacuum and clean interiors and wash and polish exteriors of automobiles. 0.925 0.828
Table 2: The two occupation profiles and two tasks from O*NET, shown in Table 1, enriched with the centroid of semantic similarity and consensus measure.
Figure 2: Heat map between cosine similarity of textual motivation of LLMs and consensus measure between scores. Data are aggregated at occupation level
3.2 Step 2: Compute the AI exposure
To compute occupation exposure to AI, we follow (Felten et al., 2021) and aggregate the te scores at the occupation level by weighting them by task relevance ($R$), importance ($I$), and frequency ($F$) as measured by ONET. [^9] These weights capture different aspects of the tasks. More specifically. Importance: indicates the degree to which a particular descriptor is important for the occupation. Relevance refers to the proportion of incumbents who rated the task as relevant to their job. Frequency refers to the frequency of each task within the occupation, from annual to hourly. Despite providing task descriptions, ONET does not provide rating, importance and frequency of tasks for 39 occupations. We manually assigned values for them. More specifically, for each task $j$ and occupation $i$ our AI exposure score is computed as follows:
$TEAI_{i}=\frac{\sum^{n}_{j=1}TE_{ij}\cdot R_{ij}\cdot I_{ij}\cdot F_{ij}}{\sum^{n}_{j=1}\cdot R_{ij}\cdot I_{ij}\cdot F_{ij}}$ (2)
where $TE_{ij}$ identifies the metric developed in step 1 at task level, $n$ defines the number of tasks within each occupation. Each weight is scaled by its maximum to obtain equal weights. The ONET model uses different scales for Relevance (scale 1-100), Importance (scale 1-5), and Frequency (scale 1-7). We normalize the indexes to ensure equal scale across weights. Finally, the score is normalized to ensure comparability with other similar scores.
3.3 Step 3: Compute the AI sustitutability
The TEAI index provides a measure of exposure to AI which does not specifically address the issue of task substitution or replacement.
To deal with this, we leverage the information contained in LLM motivations obtained in the fist step. We start by considering the three motivations provided by the three LLMs in evaluating each task. We asked another LLM to generate a single motivation that summarizes the different motivations regarding the exposure rate for each task of each occupation. This summary provides a description of the task activities along with a description of what AI can contribute to task implementation.
We provided this summary along with the task title and the job occupation title to a “master” LLM [^10] A larger model than the one used in the first phase (72B versus 7B). by asking it, based on the input information, the level of engagement for the task by the AI, on a scale from 1 which denotes “no engagement” to 5 which denotes “complete replacement of the human being in performing the task”. The model was prompted to provide a motivation for the evaluation undertaken. In addition, since many tasks are in theory completely implementable by AI, we also asked the model to return a Boolean flag equal to 0 for fully automated tasks that do not require any human supervision and equal to 1 for tasks where at least some human supervision is required. [^11] As this time we did not use different LLMs, to limit hallucinations, in accordance with the literature (Ji et al., 2023), we also asked the model to return the elements given as input, i.e., occupation title and task title. During the evaluation phase, we put controls on the returned strings. The basic idea is that if the model hallucinated itself, it should return a different occupation or task title than the ones given as input.
In this way, we obtained a specific measure of task substitutability that we aggregated at the occupation level using eq (2), obtaining an index of task replacement by AI (TRAI ).
4 Human Evaluation of TEAI/TRAI indexes
To evaluate the outcomes provided by LLMs, we conducted a user study by recruiting 12 evaluators through Prolific (Palan and Schitter, 2018), an online service offering participants for research studies. It has been recognised as a cost-effective platform that provides high-quality data for online research, making it suitable for studies where reliability and data integrity are priorities (Douglas et al., 2023), and used in several computer-science related user evaluations, see, e.g. (De Filippo et al., 2024; Koppol et al., 2021).
Setting the user study
We sampled occupations from the O*NET database by SOC first-digit, selecting 20% of occupations and choosing two tasks for each occupation: one with the highest and one with the lowest TEAI/TRAI indexes. This resulted in a total of 200 occupations and 400 tasks. The sample was further divided into four chunks of 100 tasks each. Each sample was composed of job occupation title, task title, TEAI value, TEAI natural language motivation, TRAI value and TRAI natural language motivations. [^12] To avoid potential biases in human judgment, the evaluators were not informed that AI systems had generated the indexes and summaries they reviewed Each chunk was assigned to exactly three participants, whom we asked to consider each task for each job occupation and evaluate the TEAI and TRAI , with a nominal value [^13] I disagree with both; I agree only with TEAI; I agree only with TRAI; I agree with both indicating their degree of agreement with the indices. We aggregate the results obtaining agreement performance for TEAI and TRAI . It is important to highlight that these results are not intended to represent true accuracy, but a consistency with what our methodology has estimated. Consensus among the participants was also calculated using the equation 1 to assess the extent to which participants agreed on the assignment of evaluation values. Results are shown in Tab. 3.
Results
On average, evaluators tend to agree with (i) TEAI index 75% of the time (48%+27%); (ii) 71% of the time with TRAI (48%+23%), whilst (iii) only 2% of the time they disagree with the indicators. However, the reader may notice that the level of consensus among reviewers is lower than expected. This is largely because the topic at hand — the impact of AI on tasks and professions — is inherently subjective and influenced by human variability, convictions, and expectations, making evaluations highly personal and, at times, open to interpretation. Nonetheless, the results show that our indexes are meaningful as they effectively capture the current capabilities of state-of-the-art AI technologies in supporting and replacing tasks within various jobs. As a result, TEAI , TRAI , and the natural language explanations provide a valuable framework for individuals and organizations to develop policies that integrate AI into careers and the workforce.
Coherence by Chuck ( %) Consensus by Chuck
C1 C2 C3 C4 Mean C1 C2 C3 C4 Mean
Agree with both 35.8 53.8 32.0 65.9 48 .69 .72 .54 .72 .68
Agree more with TEAI 28.0 19.0 45.5 17.0 27 .66 .57 .67 .52 .61
Agree more with TRAI 35.1 26.9 20.4 11.2 23 .75 .76 .70 .64 .71
Disagree with both 1.1 0.4 2.2 5.9 2 .50 .27 .41 .28 .36
Table 3: Human evaluation results
5 Experimental results and outcomes in the US labor market
5.1 Benchmarking evaluation
First, we compare our AI index with other existing measures in the literature. Figure 3 shows the correlation between the TEAI index and the well-known measure developed by (Frey and Osborne, 2017), the AI exposure index by (Felten et al., 2021) and by (Webb, 2023), and the offshorability index developed by (Acemoglu and Autor, 2011). The pairwise correlation is always statistically significant at 5%. It is higher for the AIOE index, much lower for the AI Webb and the offshorability index, and negative for the Frey-Osborne index. This means our measure is broadly consistent with existing measures but captures different elements of the relationship between AI and the labor market. The negative correlation with the Frey and Osborne index can be explained by the latter being a measure of exposure to robotics and computerization and more focused on routine tasks. At the same time, generative AI is more focused on cognitive/non-routine tasks.
Figure 3: Correlation with existing exposure indexes. Each dot represents a SOC occupation.
5.2 AI and skills
Next, we explore the relationship between our TEAI index and different skills. Figure 4 displays scatterplots comparing the TEAI index with the intensity of different skill types at the occupation level derived from (Acemoglu and Autor, 2011). The graph shows the peculiar nature of AI technologies, which are positively correlated with cognitive analytical and interpersonal skills, but negatively correlated with routine manual skills and non-routine manual skills that require physical adaptability. Surprisingly, the correlation with cognitive routine skills is only weakly positive, while it is positive for non-routine manual skills requiring interpersonal adaptability.
Figure 4: Correlation with different skill intensity measures. Each dot represents a SOC occupation.
The results of the figure are purely descriptive; therefore, we add a more robust analysis by extracting from ONET the detailed skills associated with each occupation. We group the skills into 4 classes: Cognitive, Social, Problem Solving and Management, and Technical skills. We then develop a skill relevance index for each class at the occupation level by weighting each skill according to its level and importance. [^14] As provided by ONET. The skill relevance index is constructed as follows:
$SR_{ci}=\frac{\sum^{m}_{z=1}S_{zcj}\cdot L_{zcj}\cdot I_{zcj}}{\sum^{m}_{z=1}\cdot L_{zcj}\cdot I_{zcj}}$ (3)
where $z$ denotes the $m$ skills of class $c$ in each occupation $j$; $L$ and $I$ denote, respectively, the level and importance of each skill in each occupation.
We then estimate the following regression:
$TEAI_{i}=\alpha+\bm{S}^{\prime}_{i}\bm{\beta}+\gamma O_{i}+\epsilon_{i}$
where each observation is a SOC occupation ($i$), $TEAI_{i}$ is our measure of AI exposure, $\bm{S}$ is a 4x1 column vector of skill relevance for each of the skill classes described above at the occupation level, $\bm{\beta}$ is a 4x1 column vector of coefficients, $O$ defines occupation dummies and $\epsilon_{i}$ the error term. We saturate the model with more detailed dummies up to the fifth digit; therefore, the results are identified within each group variation. Table 4 shows the results. The TEAI index is positively correlated with cognitive, problem-solving and managerial skills, but negatively correlated with social skills, as expected. The relationship with technical skills is very weak and does not survive the inclusion of detailed SOC occupation dummies.
(1)) (2) (3)
Cognitive 5.6934∗∗∗ 5.3653∗∗∗ 5.3105∗∗∗
(0.8023) (0.7937) (1.1712)
Social -3.1395∗∗∗ -3.2464∗∗∗ -3.0310∗∗
(0.7239) (0.7335) (1.1295)
Prob. sol. man. 3.1883∗∗∗ 3.4810∗∗∗ 2.9411
(0.8183) (0.8129) (1.3644)
Technical 0.0259 -0.0419 -0.1486
(0.6432) (0.6527) (0.8556)
SOC FE 3d 4d 5d
R2 0.775 0.780 0.854
N 774 771 522
Table 4: OLS estimates of TE-AI index on measures of skill intensity
5.3 AI employment and wages
Finally, we explore the relationship between TEAI and labor market outcomes. We start by analyzing the size and the characteristics of workers exposed to AI technologies. First, we divide the distribution of TEAI scores into three tertiles representing high, medium, and low AI exposure. We then compute the degree of exposure of the US population using BLS employment data. Finally, we distinguish between occupational groups and by skill groups within each tertile.
(a) TEAI index by SOC group of O*NET
(b) TEAI index by skill intensity
Figure 5: Exposure to TEAI index by SOC group (Fig. 5(a)) and by skill intensity (Fig. 5(b)). US BLS employment. Values in millions of workers. Each bar represents a tertile of the TEAI score distribution.
Figures 5(a) and 5(b) show the results. Overall, in 2023, 34% of US employment is highly exposed to AI technologies, while medium and low exposure represent 32% and 34%, respectively. Our findings do not suggest a polarizing effect of AI exposure as found by (Frey and Osborne, 2017); on the contrary, AI seems to have a more balanced impact on the labor market. This is because our indicator is able to capture recent advances in AI, such as LLMs, which have affected occupational groups such as management, business, administration, and finance, as well as ICT and science, which are intensive in non-routine cognitive tasks. For example, AI technologies are increasingly being used to diagnose diseases, write reports, code, or brainstorm ideas in management and business. On the contrary, previous studies that focus more on the impact of AI on routine tasks find that these tasks and occupations are less exposed to AI.
Grouping occupations by skill intensity shows that in the group with high exposure to AI, 88% of employment is in high-skill jobs; in the group with medium exposure, 53% of employment is in medium-skill jobs while 40% is in high-skill jobs. In the group with the lowest exposure to AI, 67% are medium-skill jobs and 25% low-skill jobs. Overall, AI exposure disproportionately affects high-skill jobs, which are characterized by the competencies most affected by AI technologies. Next, we analyze the relationship between AI exposure and workers’ characteristics.
Figure 6 shows that TEAI exposure is higher for workers with a high level of education, particularly graduates and postgraduates. There is a slight increase in exposure with age, although the variation in exposure is really limited after the age of 30. Men are more exposed than women at all ages.
(a) Exposure by education
(b) Exposure by age and sex
Figure 6: AI exposure by workers’ characteristics.
Panel a) shows coefficients of regression of education categories on TEAI exposure (in percentiles). Covariates include age and sex. Estimates control for occupation(4d), industry(3d), state and year fixed effects. ACS individual weights are used. Robust standard errors are clustered at the industry level. Panel b) is a binscatter. The x-axis is the average age of workers in an industry-occupation-state observation in the 2022-18 ACS 5 years sample. Biscatter is computed considering education as a covariate. ACS individual weights are used.
Finally, we assess the relationship between AI exposure, employment, and wages. To compute the medium-term effect of AI in a flexible way, allowing for changes during the estimation period, we compute the log change in employment and wages over a 4-year rolling window from 2003 to 2023.151515The last 4Y variation is therefore 2019-2023. Therefore, for each 4Y window, we run the following regression:
$\Delta y_{i,j}=\alpha+\beta TEAI_{i}+\bm{Z}^{\prime}_{i,j}\bm{\gamma}+\delta_{i}+\eta_{j}+\epsilon_{i,j}$ (4)
where $\Delta y_{i,j}$ denotes the 4-year change in log employment and log wages in sector $j$ for occupation $i$, $\bm{Z}$ is a column vector of controls, $\delta_{i}$ denote occupation dummies, $\eta_{j}$ sector dummies and $\epsilon_{i,j}$ the error term. To control for possible endogeneity and omitted variable problems, we include as controls the initial level of employment, the initial level of wage, and wage squared. We also include detailed sector (NAICS) and occupation (SOC) fixed effects in the regression and cluster the errors at the sector level. Figure 7 shows that exposure to AI positively correlates with employment and wage growth throughout the entire period. This suggests that AI technologies complement labor and increase productivity, thereby boosting employment and wages in occupations with greater exposure to AI.
The presence of detailed controls at the industry and occupation level allows us to control for factors on the production side (changes in output across industries), on the demand side (changes in product demand across industries) and on the labor supply side (changes in employment across industries and occupations) that are unrelated to AI technologies and that could affect wages and employment. Moreover, the focus on a relatively short period isolates our results from long-term trends within industries and occupations.
Therefore, the positive relationship between employment and wages and AI exposure should be interpreted as meaning that occupations more exposed to AI have stronger employment and wage growth within the occupation and sector. Our results contrast with those obtained by (Acemoglu et al., 2022; Webb, 2023), who find a negative relationship between employment and wages. The potential reconciliation between our findings and theirs lies, on the one hand, in our construction of a different measure of AI exposure that emphasizes more recent advances in AI. On the other hand, our analysis focuses on changes over the last 20 years, while their analysis takes a longer-term perspective, focusing on changes over several decades.
(a) TEAI index and employment growth
(b) TEAI and wage growth
Figure 7: TEAI index, employment and wage growth.
This figure plots the effect of AI score on employment and wage growth. Estimates are obtained from equation 4, with rolling regression coefficients and 95% confidence intervals of 4-year windows, starting in 2003-2007 and ending in 2019-2023. The point estimate refers to TEAI score, and the dependent variables are annual percentage growth rates of employment and wages. Employment regression includes the log of the initial period of employment. Wage regression includes a log of initial period employment, log initial period wage and log initial period wage squared. All the regressions include occupation (SOC 4 digit) and sector (NAICS 3 digit) fixed effects. Robust standard errors clustered at NAICS level.
5.4 Task exposure or substitutability?
Whether AI will be an opportunity or a threat in the future will hinge on whether it complements or substitutes human labor.
The measure we have developed so far is relatively agnostic in this respect, as we cannot yet disentangle substitutability from complementarity. In other words, a high TEAI exposure measure does not necessarily imply full substitution of labor by technology, which, on the contrary, may rather complement human activities, leading to higher productivity without displacing labor.
As described in section 3 the TRAI index is specifically designed to assess task replacement.
Figure 8: Task exposure and task substitutability
This figure plots the TEAI index and the task replacement index TRAI . Each dot represents an O*NET occupation. Division lines are constructed using median values of the distribution of both indexes. An interactive visualisation and the full list of occupations and AI indexes are available at https://terminatoreconomy.com/home
(a) Management, Business, Admin, Finance & ICT Science
(b) Education, Social, Legal, Art & Health & Care
(c) Other Services vs Sales
(d) Farming & Fishing & Construction & Mining
(e) Production & Repair &Transport
Figure 9: Exposure and task substitutability by occupation groups.
This figure plots the TEAI index and the replacement index TRAI . Each panel represents a pair of occupation groups. Each dot represents an occupation. Division lines are constructed using median values.
What is the relationship between the TRAI and the TEAI index? On the one hand, the TRAI index can be considered as a sub-product of the TEAI index, as it is constructed by using the motivation summaries that LLMs produce in developing the TEAI . On the other hand, it differs because the task description that it analyzes is different (the TEAI uses ONET task description TRAI uses the summary of LMMs descriptions and motivations) and is designed to assess task substitutability, as the prompt asks LLM to specifically assess the possibility of replacing the human in task performance.
Figure 8 displays the two indexes. Both scores have been rescaled on a scale of 0-100 for better readability. Please note that the scaling does not imply an exact percentage. For example, a score of 80 does not mean that there is an 80% exposure or a replacement rate of 80%. Also, while the figure shows different jobs, our replacement and exposure measures are constructed at task level. So, even if an occupation has a high level of replacement, it does not mean that it can be completely replaced by AI. Instead, it means that AI can perform a large number of tasks in that occupation, either completely or partially. The figure shows that the two measures are highly correlated, i.e. that high values of the TEAI index are generally associated with high values of the replacement index.
Grouping occupations improves the explanatory power of the analysis. In Figure 9 we isolate different occupation groups using the same taxonomy as in Figure 5(a). For example, most ICT and science occupations display a high level of exposure and task substitutability. This means that for these occupations, AI can execute several tasks with a high degree of confidence. Human activity will still be required, and it is likely that these professionals will concentrate on more critical and high value added tasks (such as supervision, analysis, developing solutions, etc.) leaving other tasks to AI. Analogously most teachers are characterized by a high level of exposure by a low level of substitutability: they are likely to work with AI while AI will replace few tasks. Similarly, several health and care occupations are characterized by low level of exposure and substitutability with few exceptions, such as “Medical record specialists” whose tasks are highly replaceable by technology. This provides some insights about the complementarity vs substitutability debate. Complementarity is often interpreted as AI and humans sharing the same activities. When looking at specific tasks, it emerges clearly that most of complementarity operates through substitution; in other words, some tasks will be executed by AI and machines while humans will concentrate on others. AI and humans will complement each other in the same occupation while the task allocation within occupation is likely to change.
6 Conclusions
We proposed a methodology for assessing AI exposure using LLMs, with a fine-grained approach that analyses exposure for 923 jobs and their 19281 tasks, using the O*NET US official taxonomy as a baseline. Our approach differs from the literature in that we provide a data-driven evaluation, where LLM systems are asked to assess the suitability of tasks for AI, resulting in an occupation-based score of AI exposure and replacement. Our index of AI exposure, the so-called TEAI is positively correlated with cognitive, problem-solving and management skills, highlighting the role of recent advances in AI that have a strong impact on management and decision-making tasks; conversely, our measure is negatively correlated with social skills, a known weakness of AI.
Regarding labor market outcomes, we find that AI exposure positively correlates with employment and wage growth over the period 2003-2023, suggesting that AI has a positive effect on productivity. Therefore, at least in the medium term, AI positively impacts the labor market. However, our estimates show that about one-third of the American workforce is at high risk of AI exposure, most in high-skilled jobs. For these workers, the overall impact of AI depends on whether exposure is associated with task substitutability. We then construct a specific version of the AI index to assess task substitutability by AI, namely TRAI . We find that even in high-skill occupations, there is high variability in task substitution by AI, suggesting that AI and humans complement each other in the same occupation while allocating tasks within professions is likely to change.
A human evaluation has been performed to assess the approach’s effectiveness, revealing that an average of 71-75% of the time evaluators agree with the score and explanations provided by our indexes, whilst only 2% of the time tend to disagree fully. To ensure full reproducibility, our methodology relies only on top-performing open LLMs, and all occupation and task lists are accessible through an online demo. This would allow the community to track future assessments of improvements in new LLM models while integrating our findings to help individuals and organizations shape policies for AI integration into careers and the workforce.
Finally, while O*NET offers a comprehensive and widely recognized taxonomy of the labour market, its static nature may not capture emerging tasks as they emerge in the real-labour market. To address this limitation, our approach is designed to be flexible and adaptable to any user-defined taxonomy, enabling organizations to assess the impact of AI within their operational contexts accurately.
Online Demo. We have made the results of our framework shown in this paper available at https://terminatoreconomy.com/home.

### v3

@@
 Many researchers have been working on estimating if and to what extent jobs and tasks are exposed to the risk of being automatized by AI-related technologies.
-Our work tackles this issue through a data-driven approach by: (i) developing a reproducible framework that uses cutting-edge open-source large language models to assess the current capabilities of AI and robotics in performing job-related tasks; (ii) formalizing and computing a measure of AI exposure by occupation, the TEAI (Task Exposure to AI) index, and a measure of Task Replacement by AI, the TRAI index, both validated through a human user evaluation and compared with the state of the art.
-Our results show that the TEAI index is positively correlated with cognitive, problem-solving and management skills, while it is negatively correlated with social skills.
-Applying the index to the US, we obtain that about one-third of US employment is highly exposed to AI, primarily in high-skill jobs, requiring a graduate or postgraduate level of education.
-We also find that AI exposure is positively associated with both employment and wage growth in 2003-2023, suggesting that AI has an overall positive effect on productivity.
+Our work tackles this issue through a data-driven approach by: (i) developing a reproducible framework that uses cutting-edge open-source large language models to assess the current capabilities of AI and robotics in performing job-related tasks; (ii) formalising and computing a measure of AI exposure by occupation, the TEAI (Task Exposure to AI) index, and a measure of Task Replacement by AI, the TRAI index, both validated through a Human user evaluation and compared with the state-of-the-art.
+Our results show that the TEAI index is positively correlated with cognitive, problem-solving, and management skills, while it is negatively correlated with social skills.
+Results also suggest about one-third of U.S.
+employment is highly exposed to AI, primarily in high-skill jobs requiring a graduate or postgraduate level of education.
+We also find that AI exposure is positively associated with employment and wage growth in 2003–2023, suggesting that AI has had an overall positive effect on productivity.
 Considering specifically the TRAI index, we find that even in high-skill occupations, AI exhibits high variability in task substitution, suggesting that AI and humans complement each other within the same occupation, while the allocation of tasks within occupations is likely to change.
 All results, models, and code are freely available online to allow the community to reproduce our results, compare outcomes, using our work as a benchmark to monitor AI’s progress over time.
-Keywords:
-AI, Large Language models, employment, skills.
 1 Introduction
-The famous 1984 movie “The Terminator” is set in a dystopian future where intelligent machines, created by a military defense system known as Skynet, become self-aware and perceive humanity as a threat, initiating a war to eliminate humans.
-Skynet creates advanced humanoid robots called Terminators to hunt down and kill human survivors.
-The Terminator has advanced learning algorithms that enable it to adapt to any environment, making it a formidable antagonist for humans.
-The debate and concerns about the impact of AI are often framed against the backdrop of the film’s setting.
-This paper takes these concerns seriously by developing an AI-centered assessment of the potential exposure of different occupations to artificial intelligence.
-Assessing the impact of technology on the labor market is not easy, as there are several potential channels at work.
-As stressed by (Acemoglu and Restrepo, 2019), technology has three major effects on labor demand.
-The first is the productivity effect, which operates through lower production costs brought about by new technologies.
-The second is the displacement effect of workers by machines and the like.
-These two effects operate in different directions and depend on whether technology substitutes for or complements human labor; in economic jargon, they depend on the elasticity of substitution between tasks.
-In addition, there is a third effect of technology: the creation of new tasks and activities in which labor can be productively employed (the reinstatement effect).
-Indeed, if we look at history, the reinstatement effect has been a central feature of all technological revolutions, which have continually created new opportunities for labor.
-For the reinstatement effect to take place, technology must have an impact wider than its narrow scope, with spillover effects in sectors/areas other than those for which it was designed.
-In other words, technology must have the features of a general purpose technology, which, according to (Lipsey et al., 2005), are pervasiveness across the economy, ability to generate complementary innovations, and improvement over time.
-AI provides strong arguments for being considered a general-purpose technology due to its broad applicability, potential productivity gains, and possibility to drive further innovation.
+Artificial Intelligence (AI) provides strong arguments for being considered a general-purpose technology due to its broad applicability, potential productivity gains and possibility to drive further innovation Crafts-21; Brynjolfsson-Rock-Syverson-19; deming2025technological; WEF2025.
 However, these characteristics of AI create a relevant measurement problem, as it is extremely difficult to identify all the channels through which it affects the economy.
-This paper contributes to this field by developing a methodology for assessing AI exposure using Large Language Models (LLMs), with a granular approach that analyzes exposure for each task within each occupation.
-Our approach allows us to improve our understanding of the effect of AI on the labor market in several areas.
-Methodologically, we design and implement a reproducible framework to assess to what extent existing AI and robotics technologies can perform job-related tasks, relying LLMs.
-Rather than assessing AI exposure through external benchmarks such as expert judgment or AI patents and innovations data, we construct an internal assessment using LLM’s own evaluation.
-In other words, LLMs assess their ability to perform each of the 19K job-related tasks associated with about 1k occupations as described by O*NET [^1] O*NET: Occupational Information Network, is a comprehensive database containing detailed information on hundreds of standardized and occupation-specific descriptors.
-It is sponsored by the U.S.
-Department of Labor/Employment and Training Administration.
-O*NET serves as a resource to provide information on skills, abilities, knowledge, work activities, and interests associated with occupations.
-This approach has two main advantages.
-On the one hand, it is very granular.
-Each individual task, in fact, receives a specific assessment and a motivation by each LLM.
-On the other hand, it is fully transparent.
-Not only does it allow full reproducibility, [^2] All codes are available on GitHub at https://github.com/Crisp-Unimib/Terminator-Economy but more importantly, it allows comparability over time.
-AI is an innovative and fast-growing field where new models are expected to significantly outperform existing ones.
-New LLMs can be applied to our approach to measure the extent of their improvement in task performance.
-Consider taxi drivers, often used as an example of a potentially exposed occupation.
-The current assessment of the task of driving vehicles by LLMs is of average exposure, recognizing that current technologies can replace some tasks but not others that need to be performed by humans.
-However, with advances in AI systems, it is likely that the same activity will be fully automated in a few years’ time.
-New models will therefore assign a higher score and allow the extent of the change in exposure to be measured.
-From an economic perspective, we develop an AI exposure measure for each O*NET occupation based on specific assessment of individual tasks .
-Our AI exposure index is positively correlated with cognitive, problem-solving and management skills, consistent with the evidence that AI affects management and decision-making tasks; on the other hand, AI exposure is negatively correlated with social skills, where AI has clear limitations.
-We also develop a version of the index which is specifically constructed for measuring task substitutability by AI and which can be used to analyze the relationship between AI exposure and AI task substitutability.
-This analysis is enhanced by the fact that our approach provides, for each task, a description of AI involvement which constitutes a valuable wealth of information that can be used to analyze the human-machine interaction.
-Bringing our measure to the data, we show that, in the US, approximately 1/3 of employment is highly exposed to AI technologies; most of this employment (88%) is in high-skill jobs, requiring graduate or postgraduate level of education.
-In terms of labor market outcomes, using 4-year rolling regressions, we find that AI exposure is positively associated with both employment and wage growth in the period 2003-2023, suggesting that AI has an overall positive effect on productivity.
-The remainder of the paper is structured as follows: section 2 discusses the related literature, section 3 describes the methodology and the construction of the AI index, section 5 presents the results; finally, section 6 concludes.
-2 Background and Related Works
-AI and jobs
-Since the seminal paper by (Autor et al., 2003), the task approach has proved very effective in analyzing the impact of technology on jobs.
-It divides work activities into tasks, each of which can be performed by humans or by machines.
-In this way, the distinction between capital and labor tasks is more precise, flexible, and able to shift over time.
-Indeed, capital and machines can substitute for labor in the performance of some tasks, while complementing it in the performance of others.
-The task approach has been applied to analyze the effect of technology and trade (offshoring) (Acemoglu and Autor, 2011), to the long run effect of technology (Consoli et al., 2023) and to skill-task interaction (Colombo et al., 2019).
-More recently this approach has been used to measure occupational exposure to computers and robots.
-In a seminal paper, (Frey and Osborne, 2017) estimated that up to 47% of jobs in the US are at risk of automation.
-[^3] See also (Nedelkoska and Quintini, 2018) for a similar approach.
-Subsequently, other attempts focused on developing measures of exposure to machine learning and robotics (Brynjolfsson and Mitchell, 2017; Acemoglu and Restrepo, 2020) and to AI (Felten et al., 2021; Webb, 2023; Eloundou et al., 2024; Pizzinelli et al., 2023).
-Overall, these works find an extensive share of employment exposed to AI; the specific impact on occupations varies depending on the nuances that different indicators of the impact of technology capture, i.e., whether they focus on aspects of technology that impact more routine-based activities (Frey and Osborne, 2017) or more cognitive elements (Felten et al., 2021).
-A common feature of these papers is the attempt to quantify AI exposure through an external benchmark, which may be expert judgment or data analysis on patents and innovations.
+Meanwhile, in late 2018, Large Language Models (LLMs) were introduced as powerful computational tools designed to understand and generate human-like text by leveraging vast amounts of data, see min2023recent.
+Due to their performance and adaptability as off-the-shelf models, they have become central to numerous systems capable of processing and generating text, driving applications in natural language understanding, translation, and content creation across domains.
+Contribution.
+This paper develops a methodology for assessing AI exposure using LLMs, with a granular approach that analyses exposure for each job task within each job occupation from two different perspectives:
+(i) From a methodological point of view, we design and implement a reproducible framework to estimate to what extent existing AI and robotics technologies can perform job-related tasks relying on LLMs.
+Instead of assessing AI exposure through external benchmarks such as expert judgment or AI patents and innovations data, we construct an internal assessment using LLM’s own evaluation, using O*NET [^1] O*NET is the official US comprehensive taxonomy of about 1K occupations and 19K+ job-related tasks as reference occupation/task taxonomy.
+The result is a framework that enables individuals and organizations to develop policies for integrating AI into careers and the workforce.
+(ii) From an economic perspective, we develop a measure to estimate the AI exposure and replacement index by tasks for all the occupations of the US taxonomy.
+We conduct an in-depth analysis of the impact of AI across occupations in the US labor market, examining job characteristics such as skill levels and educational requirements.
+We investigate the relationship between AI exposure and labor market outcomes like employment and wage growth over 2019–2023.
+This combined analysis allows us to estimate the impact of AI and technology on different occupations and economic indicators.
+Reproducibility and Demo.
+Codes and results are available on GitHub [^2] https://github.com/Crisp-Unimib/Terminator-Economy while an interactive online demo is available at http://terminatoreconomy.com/
+2 Related Works on AI and Jobs
+Since the seminal paper by Autor-Levy_Murnane-2003, the task approach has been applied to analyze the effect of technology and trade (offshoring) Acemoglu-Autor-11, to the long run effect of technology Consoli-Marin-Rentocchini-Vona-2023, and to skill-task interaction Colombo-Mercorio-Mezzanzanica-2019.
+More recently, this approach has been used to measure occupational exposure to computers and robots.
+In a seminal paper Frey-Osborne-17 estimated that up to 47% of jobs in the US are at risk of automation.
+[^3] See also Nedelkoska-Quintini-18 for a similar approach.
+Subsequently, other attempts focused on developing measures of exposure to machine learning and robotics Brynjolfsson-Mitchell-17; Acemoglu-Restepo-2020 and to AI Felten-Raj-Seamans-21; Webb-23; Eloundou-Manning-Mishkin-Rock-23; Pizzinelli-Panton-Tavares-Cazzaniga-Li-23.
+A common feature of those is the attempt to quantify AI exposure through an external benchmark, which may be expert judgment or data analysis on patents and innovations.
 In contrast, our approach is based on an internal assessment, where LLM systems are asked to assess the suitability of tasks for AI.
@@
 Second, the approach is fully reproducible.
-This means that whenever new generations of LLM are available, they can be used in our approach to measure the change in task exposure that they imply.
-Large Language Models
-LLMs are powerful computational models designed to understand and generate human-like text by harnessing vast amounts of textual data, and have taken Natural Language Processing (NLP) by storm, achieving state-of-the-art performance in many tasks (Min et al., 2023).
-Typically, these models are based on Transformer architecture (Vaswani et al., 2017), powered by Attention mechanism (Luong et al., 2015; Bahdanau et al., 2014) and are composed by decoder-only stack.
-These models are initially trained on an Autoregressive task Radford et al.
-(2018), where given a sequence of words $S=(w_{1},w_{2},...,w_{n-1})$ the training objective is to maximize the log-likelihood
-$\sum_{i}\log P(w_{i}|w_{1},w_{2},....,w_{i-1};\theta^{T})$ where $\theta^{T}$ are the model parameters, in order to predict the next word in the sequence $\prod_{i=1}^{n}P(w_{i}\mid w_{1},...,w_{i-1})$.
-After being pre-trained, these models are fine-tuned for several tasks, providing examples of Natural Language Inference (Radford et al., 2018).
-Thanks to their capability to learn from context, known as in-context learning (Radford et al., 2019), LLMs can accomplish specific tasks with high accuracy (Zhao et al., 2023), exploiting prompt engineering methodologies such as zero-shot (Wei et al., 2021) and few-shot learning (Brown et al., 2020).
+This means that whenever new generations of LLMs are available, they can be used in our approach to measure the change in task exposure that they imply.
 In February 2025, Anthropic published a study [^4] https://www.anthropic.com/news/the-anthropic-economic-index to assess AI’s impact on jobs and tasks using O*NET, mapping millions of Claude conversations to O*NET tasks.
 By contrast, our approach (i) maps every O*NET occupation and task, (ii) employs a panel of three open LLMs to ensure reproducibility and minimize hallucinations, and (iii) allows the community to estimate the AI impact on other third-party taxonomies.
-3 Building the AI Exposure Index
-Our method can be summarized as follows: First, we obtain from O*NET the description of each task associated with each SOC occupation.
-Second, we apply LLMs to the task descriptions to obtain a rating of how well AI technologies can perform each task.
-Third, we aggregate the rating at the occupation level to obtain an AI exposure score.
-Finally, we apply our score to US data to assess the extent of AI exposure in the US labor market and the impact of AI on employment and wages.
-Figure 1 provides a graphical representation of our approach.
-Figure 1: Graphical overview of the framework to compute the TEAI Index
+3 Building the AI Exposure and Substitutability Indexes
+Our method can be summarised as in Fig. 1.
+Figure 1: Graphical overview of the framework
 3.1 Step 1: Compute the AI Rate
-In order to limit the risk of being driven by the well-known problem of LLMs called "hallucinations" (Ji et al., 2023), we design a framework involving three different LLMs aimed at identifying and limiting the false information generated, creating a consensus system between them.
-Model choice
-To ensure the reproducibility of this work we use three of the best-performing open source models, according to performance benchmarks, available on the open LLM rankings.
-[^5] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard The three selected models are Mistral 7B Instruct v 0.2 (Jiang et al., 2023), openchat 3.5 0106 (Wang et al., 2023) and orca mini v3 7b (Mukherjee et al., 2023).
-Prompt design
+To mitigate the risk of being driven by the well-known problem of LLMs, known as ”hallucinations” ji2023survey, we design a framework involving three different LLMs that aim to identify and limit the false information generated, thereby creating a consensus system among them.
+Model choice.
+To ensure the reproducibility of this work, we use three of the best-performing open-source models (at the time of the experimental phase), according to performance benchmarks, available in the open LLM ranking.
+[^5] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard The three selected models are Mistral 7B Instruct v 0.2 jiang2023mistral, openchat 3.5 0106 wang2023openchat and orca mini v3 7b mukherjee2023orca.
+Prompt design.
 The starting point is the ONET taxonomy which identifies 19281 tasks for 923 SOC occupations.
 [^6] We use the ONET 28.2 version released in February 2024.
-We formulate a five-shot prompt using the few-shot learning approach (Brown et al., 2020).
+We formulate a five-shot prompt using the few-shot learning approach brown2020language.
 We use each individual task description assigned to an occupation to ask the models how well, on a scale of 1 to 5, [^7] How well an AI system, which can be an LLM, Image Processing System or a Robot, could perform in the task on a scale of 1 to 5 where 1 stands for poor and 5 stands for excellent?
@@
 At the end of the prompt, the model is provided with five examples of the task to obtain more contextual and accurate results.
-Model results
-We iterate this process for each task provided by ONET and for each model chosen, ending up with three scores and natural language motivations provided by each model.
+Model results.
+We iterate this process for each task provided by ONET and for each model chosen, resulting in three scores and natural language motivations provided by each model.
 Table 1 provides an example of the results after this stage for a selection of occupations and tasks.
-O*NET Task Mistral Orca-mini Open-Chat Motivation provided by Mistral Direct or coordinate the supportive services department of a business, agency, or organization.
-4 4 3 Robotics can automate repetitive tasks such as data entry, document scanning, and mail sorting.
-Image Processing Systems can analyze documents and emails for information extraction and categorization.
-LLMs can handle complex tasks like scheduling, coordinating resources, and communicating with team members.
-Combining these AI technologies can significantly improve the efficiency and accuracy of administrative services, allowing managers to focus on more strategic tasks.
-Conduct classes to teach procedures to staff.
-2 3 2 LLMs could potentially create instructional materials, but they lack the ability to engage and assess staff performance during classes.
-Image Processing Systems have no role in this task.
-Robotics could theoretically conduct classes through pre-recorded videos, but the human interaction and customization required for effective training make this a poor solution.
+O*NET Task Mistral
+Orca
+mini
+Open
+Chat
+Motivation
+Mistral
+Direct or
+coordinate
+the
+supportive
+services
+department
+of
+a business
+agency, or
+organization.
+4 4 3
+Robotics can automate
+repetitive tasks such as
+data entry, document scanning,
+and mail sorting. Image
+Processing Systems can
+analyze documents and emails
+for information extraction and
+categorization. LLMs can handle
+complex tasks like scheduling,
+coordinating resources, and
+communicating with team
+members. Combining these
+AI technologies can significantly
+improve the efficiency and
+accuracy of administrative
+services, allowing managers to
+focus on more strategic tasks.
+Conduct
+classes to
+teach
+procedures
+to staff.
+2 3 2
+LLMs could potentially
+create instructional materials,
+but they lack the ability to
+engage and assess staff
+performance during classes.
+Image Processing Systems
+have no role in this task.
+Robotics could theoretically
+conduct classes through
+pre-recorded videos, but
+the human interaction and
+customization required for
+effective training make this a
+poor solution.
 Table 1: Extraction of two tasks for Administrative Services Managers occupation O*NET, enriched as discussed in Step1 to derive to what extent existing AI and robotics technologies can perform job-related tasks according to Mistral, Orca-mini, and Open-Chat LLMs.
@@
 Key activities are highlighted for a better reading
-Evaluating consensus among LLMs
+Evaluating Consensus among LLMs.
 As mentioned above, the decision to use three different models was made to limit hallucinations.
-In order to construct a single indicator, we took a conservative approach by assigning to each task the value of the rating with the highest frequency among the three models; if the three rates were different, we chose the lowest.
-To assess the agreement between the three rates expressed by the LLMs, for each single task we compute a consensus metric (Tastle and Wierman, 2007).
-$Cns(a_{i})=1+\sum_{k=1}^{m}p_{k}\log_{2}(1-\frac{\lvert LV_{k}-\mu_{LV}\rvert}{d_{LV}})$ (1)
-Equation (1) shows the consensus calculation in which, given an input task $a_{i}$, for which three rates are expressed by three different LLMs, $LV_{k}$ denotes the rate (on a scale from 1 to 5) expressed by the $k$-th LLM, $p_{k}$ denotes the relative frequency associated with $LV_{k}$; $\mu_{LV}$ defines the weighted average of the three $LV$ rating values computed using $p_{k}$ probabilities as weights; finally $d_{LV}$ represents the range of the rating scale adopted.
-Specifically, for each $k$-th rate from 1 to 3 (corresponding to the three available rates), the absolute difference between the rating value $LV_{k}$ and the weighted average $\mu_{LV}$ is determined.
-This difference is then normalized by dividing by the scale size $d_{LV}$.
-The logarithmic function $\log_{2}$ is applied in order to moderate the impact of larger discrepancies from the mean.
-This process allows the three rates to be combined into a single measure of agreement.
-A higher consensus value indicates greater agreement between the observed rating values expressed for each task by the three different LLMs.
-Analogously, to estimate the similarity between the motivations provided by the LLMs, we compute the centroid of semantic cosine similarity (Rahutomo et al., 2012), between them.
-The embedding vectors for the centroid computation are obtained using an open source Transformer model: as for the LLMs, we chose the Transformer model to be used in accordance with the Massive Text Embedding Benchmark (MTEB) Leaderboard by (Muennighoff et al., 2023).
-[^8] https://huggingface.co/spaces/mteb/leaderboard.
-Having English-language motivations, the choice fell on the UAE-Large-V1 https://huggingface.co/WhereIsAI/UAE-Large-V1, which represented an excellent compromise between effectiveness and efficiency, given its small size.
-Table 2 shows the results following the calculation of semantic similarity and consensus between the different rates.
-In particular, since higher cosine similarity values reflect higher semantic similarities between the text of the LLM motivations, we expect a strong correlation between the consensus metric and cosine similarity.
-The heat map represented in figure 2 shows that both the values of cosine similarity and the consensus metric are extremely high, with an average close to 0.9 in both cases.
-On the one hand, this suggests coherence between LLM-generated rates and the associated motivations, and on the other hand, it adds robustness to our conservative approach in choosing between different models.
-The result of this process is a single score te that returns a metric from 1 to 5 measuring the extent to which AI can perform each specific task, and a quantitative indicator of the similarity between the discursive motivations generated by the models.
-O*NET Occupation O*NET Task Similarity Consensus
-Administrative Services Managers Direct or coordinate the supportive services department of a business, agency, or organization. 0.918 0.828
-Conduct classes to teach procedures to staff. 0.935 0.828
-Taxi Drivers Complete accident reports when necessary. 0.948 0.828
-Vacuum and clean interiors and wash and polish exteriors of automobiles. 0.925 0.828
-Table 2: The two occupation profiles and two tasks from O*NET, shown in Table 1, enriched with the centroid of semantic similarity and consensus measure.
-Figure 2: Heat map between cosine similarity of textual motivation of LLMs and consensus measure between scores. Data are aggregated at occupation level
+In Section A.1, we describe the analysis conducted on the output of the three models and how we obtain a single indicator for each task, defined as te, a metric ranging from 1 to 5 that measures the extent to which AI can perform each specific task.
 3.2 Step 2: Compute the AI Exposure
-To compute occupation exposure to AI, we follow (Felten et al., 2021) and aggregate the te scores at the occupation level by weighting them by task relevance ($R$), importance ($I$), and frequency ($F$) as measured by ONET.
-[^9] These weights capture different aspects of the tasks.
-More specifically.
+To compute occupation exposure to AI, we follow Felten-Raj-Seamans-21 and aggregate the te scores at the occupation level by weighting them by task relevance ($R$), importance ($I$), and frequency ($F$) as measured by ONET.
+[^8] Weights capture different aspects of the tasks.
 Importance: indicates the degree to which a particular descriptor is important for the occupation.
@@
 Frequency refers to the frequency of each task within the occupation, from annual to hourly.
-Despite providing task descriptions, ONET does not provide rating, importance and frequency of tasks for 39 occupations.
-We manually assigned values for them.
-More specifically, for each task $j$ and occupation $i$ our AI exposure score is computed as follows:
+For each task $j$ and occupation $i$ our AI exposure score is computed as follows:
 $TEAI_{i}=\frac{\sum^{n}_{j=1}TE_{ij}\cdot R_{ij}\cdot I_{ij}\cdot F_{ij}}{\sum^{n}_{j=1}\cdot R_{ij}\cdot I_{ij}\cdot F_{ij}}$ (1)
-where $TE_{ij}$ identifies the metric developed in step 1 at task level, $n$ defines the number of tasks within each occupation.
+where $TE_{ij}$ identifies the metric developed in step 1 at the task level, $n$ defines the number of tasks within each occupation.
 Each weight is scaled by its maximum to obtain equal weights.
 The ONET model uses different scales for Relevance (scale 1-100), Importance (scale 1-5), and Frequency (scale 1-7).
-We normalize the indexes to ensure equal scale across weights.
+We normalize the indices to ensure equal scale across weights.
 Finally, the score is normalized to ensure comparability with other similar scores.
 3.3 Step 3: Compute the AI Sustitutability
-The TEAI index provides a measure of exposure to AI which does not specifically address the issue of task substitution or replacement.
-To deal with this, we leverage the information contained in LLM motivations obtained in the fist step.
-We start by considering the three motivations provided by the three LLMs in evaluating each task.
-We asked another LLM to generate a single motivation that summarizes the different motivations regarding the exposure rate for each task of each occupation.
-This summary provides a description of the task activities along with a description of what AI can contribute to task implementation.
-We provided this summary along with the task title and the job occupation title to a “master” LLM [^10] A larger model than the one used in the first phase (72B versus 7B).
-by asking it, based on the input information, the level of engagement for the task by the AI, on a scale from 1 which denotes “no engagement” to 5 which denotes “complete replacement of the human being in performing the task”.
-The model was prompted to provide a motivation for the evaluation undertaken.
-In addition, since many tasks are in theory completely implementable by AI, we also asked the model to return a Boolean flag equal to 0 for fully automated tasks that do not require any human supervision and equal to 1 for tasks where at least some human supervision is required.
-[^11] As this time we did not use different LLMs, to limit hallucinations, in accordance with the literature (Ji et al., 2023), we also asked the model to return the elements given as input, i.e., occupation title and task title.
-During the evaluation phase, we put controls on the returned strings.
-The basic idea is that if the model hallucinated itself, it should return a different occupation or task title than the ones given as input.
-In this way, we obtained a specific measure of task substitutability that we aggregated at the occupation level using eq (2), obtaining an index of task replacement by AI (TRAI ).
+The TEAI index measures exposure to AI but does not specifically address the issue of task substitution or replacement.
+To deal with this, we leverage information contained in the motivations obtained in the first step.
+We synthesize these, generating a consolidated motivation that summarizes the different justifications regarding the AI exposure rate for each occupation-related task.
+To obtain a quantitative measure of AI’s involvement in each task, we provide this summarized motivation along with the task title and the associated job occupation title to another LLM,prompting it to return the level of AI engagement on a scale from 1 (“no engagement”) to 5 (“complete automation, replacing human execution”) and the motivation regarding the engagement.
+Model choice and Results.
+In this step we used a single LLM that was larger than those used in the first step.
+To ensure the complete reproduction of the results, we chose the Qwen 72B model, which, among open-source models, represents a good trade-off between size and performance.
+The result is a measure of task substitutability that we aggregated at the occupation level using eq (1), obtaining an index of Task Replacement by AI (TRAI).
 4 Human Evaluation of TEAI/TRAI Indexes
-To evaluate the outcomes provided by LLMs, we conducted a user study by recruiting 12 evaluators through Prolific (Palan and Schitter, 2018), an online service offering participants for research studies.
-It has been recognised as a cost-effective platform that provides high-quality data for online research, making it suitable for studies where reliability and data integrity are priorities (Douglas et al., 2023), and used in several computer-science related user evaluations, see, e.g.
-(De Filippo et al., 2024; Koppol et al., 2021).
-Setting the user study
+To evaluate the outcomes provided by LLMs, we conducted a user study by recruiting 12 evaluators through Prolific palan2018prolific, an online service offering participants for research studies.
+It has been recognized as a cost-effective platform that provides high-quality data for online research, making it suitable for studies where reliability and data integrity are priorities douglas2023data, and used in several computer-science related user evaluations, see, e.g., de2024large; koppol2021interaction.
+Setting the user study.
 We sampled occupations from the O*NET database by SOC first-digit, selecting 20% of occupations and choosing two tasks for each occupation: one with the highest and one with the lowest TEAI/TRAI indexes.
@@
 The sample was further divided into four chunks of 100 tasks each.
-Each sample was composed of job occupation title, task title, TEAI value, TEAI natural language motivation, TRAI value and TRAI natural language motivations.
+Each sample was composed of job occupation title, task title, TEAI value, TEAI natural language motivation, TRAI value, and TRAI natural language motivations.
 [^9] To avoid potential biases in human judgment, the evaluators were not informed that AI systems had generated the indexes and summaries they reviewed Each chunk was assigned to exactly three participants, whom we asked to consider each task for each job occupation and evaluate the TEAI and TRAI , with a nominal value [^10] I disagree with both; I agree only with TEAI; I agree only with TRAI; I agree with both indicating their degree of agreement with the indices.
-We aggregate the results obtaining agreement performance for TEAI and TRAI .
+We aggregate the results, obtaining agreement performance for TEAI and TRAI.
 It is important to highlight that these results are not intended to represent true accuracy, but a consistency with what our methodology has estimated.
@@
 Results are shown in Tab.
-3.
-Results
-On average, evaluators tend to agree with (i) TEAI index 75% of the time (48%+27%); (ii) 71% of the time with TRAI (48%+23%), whilst (iii) only 2% of the time they disagree with the indicators.
-However, the reader may notice that the level of consensus among reviewers is lower than expected.
+2.
+Results.
+Averagely, evaluators tend to agree with (i) TEAI index 75% of the time (48%+27%); (ii) 71% of the times with TRAI (48%+23%), whilst (iii) only 2% of the times they disagree with the indicators.
+However, the reader may notice that the level of consensus among reviewer is lower than expected.
 This is largely because the topic at hand — the impact of AI on tasks and professions — is inherently subjective and influenced by human variability, convictions, and expectations, making evaluations highly personal and, at times, open to interpretation.
-Nonetheless, the results show that our indexes are meaningful as they effectively capture the current capabilities of state-of-the-art AI technologies in supporting and replacing tasks within various jobs.
+Nonetheless, the results show that our indices are meaningful as they effectively capture the current capabilities of state-of-the-art AI technologies in supporting and replacing tasks within various jobs.
 As a result, TEAI, TRAI, and the natural language explanations provide a valuable framework for individuals and organizations to develop policies that integrate AI into careers and the workforce.
@@
 C1 C2 C3 C4 Mean C1 C2 C3 C4 Mean
-Agree with both 35.8 53.8 32.0 65.9 48 .69 .72 .54 .72 .68
-Agree more with TEAI 28.0 19.0 45.5 17.0 27 .66 .57 .67 .52 .61
-Agree more with TRAI 35.1 26.9 20.4 11.2 23 .75 .76 .70 .64 .71
-Disagree with both 1.1 0.4 2.2 5.9 2 .50 .27 .41 .28 .36
-Table 3: Human evaluation results
-5 Experimental results and outcomes in the US labor market
-5.1 Benchmarking evaluation
+Agree with both 35.84 53.76 32 65.94 48 .69 .72 .54 .72 .68
+Agree more with TEAI 27.96 19 45.48 17.03 27 .66 .57 .67 .52 .61
+Agree more with TRAI 35.13 26.88 20.35 11.23 23 .75 .76 .70 .64 .71
+Disagree with both 1.08 0.36 2.17 5.89 2 .50 .27 .41 .28 .36
+Table 2: Human evaluation results.
+Mean of Coherence and Consensus Levels among reviewers by chunk.
+Each chunk has been assigned to three reviewers.
+The closer the consensus is to 1, the more the different reviewers agree.
+5 Experimental Results and Outcomes on the US Labor Market
+Benchmarking evaluation.
 First, we compare our AI index with other existing measures in the literature.
-Figure 3 shows the correlation between the TEAI index and the well-known measure developed by (Frey and Osborne, 2017), the AI exposure index by (Felten et al., 2021) and by (Webb, 2023), and the offshorability index developed by (Acemoglu and Autor, 2011).
+Figure 2 shows the correlation between the TEAI index and the well-known measure developed by Frey-Osborne-17, the AI exposure index by Felten-Raj-Seamans-21 and by Webb-23, and the offshorability index developed by Acemoglu-Autor-11.
 The pairwise correlation is always statistically significant at 5%.
@@
 This means our measure is broadly consistent with existing measures but captures different elements of the relationship between AI and the labor market.
-The negative correlation with the Frey and Osborne index can be explained by the latter being a measure of exposure to robotics and computerization and more focused on routine tasks.
+The negative correlation with the Frey and Osborne index can be explained by the latter being a measure of exposure to robotics and computerisation, and more focused on routine tasks.
 At the same time, generative AI is more focused on cognitive/non-routine tasks.
 Figure 2: Correlation with existing exposure indexes. Each dot represents a SOC occupation.
+Figure 3: Correlation with different skill intensity measures.
 5.1 AI and Skills
 Next, we explore the relationship between our TEAI index and different skills.
-Figure 4 displays scatterplots comparing the TEAI index with the intensity of different skill types at the occupation level derived from (Acemoglu and Autor, 2011).
-The graph shows the peculiar nature of AI technologies, which are positively correlated with cognitive analytical and interpersonal skills, but negatively correlated with routine manual skills and non-routine manual skills that require physical adaptability.
+Figure 3 displays scatterplots comparing the TEAI index with the intensity of different skill types at occupation level derived from Acemoglu-Autor-11.
+The graph shows the peculiar nature of AI technologies, which are positively correlated with cognitive analytical and interpersonal skills but negatively correlated with routine manual skills and non-routine manual skills that require physical adaptability.
 Surprisingly, the correlation with cognitive routine skills is only weakly positive, while it is positive for non-routine manual skills requiring interpersonal adaptability.
-Figure 4: Correlation with different skill intensity measures. Each dot represents a SOC occupation.
-The results of the figure are purely descriptive; therefore, we add a more robust analysis by extracting from ONET the detailed skills associated with each occupation.
+The results of the figure are purely descriptive, therefore we add a more robust analysis by extracting from ONET the detailed skills associated with each occupation.
 We group the skills into 4 classes: Cognitive, Social, Problem Solving and Management, and Technical skills.
@@
 Table 3 shows the results.
-The TEAI index is positively correlated with cognitive, problem-solving and managerial skills, but negatively correlated with social skills, as expected.
+The TEAI index is positively correlated with cognitive, problem-solving, and managerial skills, but negatively correlated with social skills, as expected.
 The relationship with technical skills is very weak and does not survive the inclusion of detailed SOC occupation dummies.
@@
 (0.7239) (0.7335) (1.1295)
-Prob. sol. man. 3.1883∗∗∗ 3.4810∗∗∗ 2.9411
+Management 3.1883∗∗∗ 3.4810∗∗∗ 2.9411
 (0.8183) (0.8129) (1.3644)
@@
 N 774 771 522
-Table 4: OLS estimates of TE-AI index on measures of skill intensity
+Table 3: OLS estimates of TEAI index on measures of skill intensity.
+Note: Each observation consists of an occupation.
+OLS regression using TEAI index as the dependent variable.
+The independent variables are skill intensities.
+All the regressions include occupation (SOC) fixed effects at 3, 4, and 5 digits.
+Robust standard errors in parentheses *** p << 0.001, ** p << 0.01, * p << 0.05
 5.2 AI Employment and Wages
 Finally, we explore the relationship between TEAI and labor market outcomes.
-We start by analyzing the size and the characteristics of workers exposed to AI technologies.
+We begin by analyzing the size and characteristics of workers exposed to AI technologies.
 First, we divide the distribution of TEAI scores into three tertiles representing high, medium, and low AI exposure.
@@
 (b) TEAI index by skill intensity
-Figure 5: Exposure to TEAI index by SOC group (Fig.
-5(a)) and by skill intensity (Fig.
-5(b)).
-US BLS employment.
-Values in millions of workers.
-Each bar represents a tertile of the TEAI score distribution.
-Figures 5(a) and 5(b) show the results.
+Figure 4: Figures 4(a) and 4(b) display the distribution of exposure to the TEAI index across SOC groups and skill intensity levels, respectively, using U.S.
+BLS employment data (in millions of workers).
+For both panels, each bar represents a tertile of the TEAI score distribution.
+Figures 4(a) and 4(b) show the results.
 Overall, in 2023, 34% of US employment is highly exposed to AI technologies, while medium and low exposure represent 32% and 34%, respectively.
-Our findings do not suggest a polarizing effect of AI exposure as found by (Frey and Osborne, 2017); on the contrary, AI seems to have a more balanced impact on the labor market.
-This is because our indicator is able to capture recent advances in AI, such as LLMs, which have affected occupational groups such as management, business, administration, and finance, as well as ICT and science, which are intensive in non-routine cognitive tasks.
+Our findings do not suggest a polarizing effect of AI exposure as found by Frey-Osborne-17; on the contrary, AI seems to have a more balanced impact on the labor market.
+This is because our indicator can capture recent advances in AI, such as LLMs, which have affected occupational groups such as management, business, administration, and finance, as well as ICT and science, which are intensive in non-routine cognitive tasks.
 For example, AI technologies are increasingly being used to diagnose diseases, write reports, code, or brainstorm ideas in management and business.
 On the contrary, previous studies that focus more on the impact of AI on routine tasks find that these tasks and occupations are less exposed to AI.
-Grouping occupations by skill intensity shows that in the group with high exposure to AI, 88% of employment is in high-skill jobs; in the group with medium exposure, 53% of employment is in medium-skill jobs while 40% is in high-skill jobs.
+Grouping occupations by skill intensity shows that in the group with high exposure to AI, 88% of employment is in high-skill jobs; in the group with medium exposure, 53% of employment is in medium-skill jobs, while 40% is in high-skill jobs.
 In the group with the lowest exposure to AI, 67% are medium-skill jobs and 25% low-skill jobs.
-Overall, AI exposure disproportionately affects high-skill jobs, which are characterized by the competencies most affected by AI technologies.
+Overall, AI exposure disproportionately affects high-skill jobs, which are characterized by the competencies most heavily affected by AI technologies.
 Next, we analyze the relationship between AI exposure and workers’ characteristics.
 Figure 5 shows that TEAI exposure is higher for workers with a high level of education, particularly graduates and postgraduates.
-There is a slight increase in exposure with age, although the variation in exposure is really limited after the age of 30.
+There is a slight increase in exposure with age, although the variation in exposure is limited after age 30.
 Men are more exposed than women at all ages.
-(a) Exposure by education
-(b) Exposure by age and sex
-Figure 6: AI exposure by workers’ characteristics.
-Panel a) shows coefficients of regression of education categories on TEAI exposure (in percentiles).
-Covariates include age and sex.
-Estimates control for occupation(4d), industry(3d), state and year fixed effects.
-ACS individual weights are used.
-Robust standard errors are clustered at the industry level.
-Panel b) is a binscatter.
-The x-axis is the average age of workers in an industry-occupation-state observation in the 2022-18 ACS 5 years sample.
-Biscatter is computed considering education as a covariate.
-ACS individual weights are used.
 Finally, we assess the relationship between AI exposure, employment, and wages.
-To compute the medium-term effect of AI in a flexible way, allowing for changes during the estimation period, we compute the log change in employment and wages over a 4-year rolling window from 2003 to 2023.151515The last 4Y variation is therefore 2019-2023.
-Therefore, for each 4Y window, we run the following regression:
+To compute the medium-term effect of AI in a flexible way, allowing for changes during the estimation period, we compute the log change in employment and wages over a 4-year rolling window from 2003 to 2023.121212The last 4Y variation is therefore 2019-2023.
+Therefore, for each 4Y window, we run the following:
 $\Delta y_{i,j}=\alpha+\beta TEAI_{i}+\bm{Z}^{\prime}_{i,j}\bm{\gamma}+\delta_{i}+\eta_{j}+\epsilon_{i,j}$ (3)
@@
 Moreover, the focus on a relatively short period isolates our results from long-term trends within industries and occupations.
-Therefore, the positive relationship between employment and wages and AI exposure should be interpreted as meaning that occupations more exposed to AI have stronger employment and wage growth within the occupation and sector.
-Our results contrast with those obtained by (Acemoglu et al., 2022; Webb, 2023), who find a negative relationship between employment and wages.
-The potential reconciliation between our findings and theirs lies, on the one hand, in our construction of a different measure of AI exposure that emphasizes more recent advances in AI.
-On the other hand, our analysis focuses on changes over the last 20 years, while their analysis takes a longer-term perspective, focusing on changes over several decades.
+Therefore, the positive relationship between employment and wages and AI exposure should mean that occupations more exposed to AI have stronger employment and wage growth within the occupation and sector.
+Our results contrast with those obtained by Acemoglu-Autor_Hazell-Restrepo-22, and Webb-23, who find a negative relationship between employment and wages.
+The potential reconciliation between our findings and theirs lies, on the one hand, in our construction of a different measure of AI exposure that emphasises more recent advances in AI.
+On the other hand, our analysis focuses on changes over the last 20 years, whereas theirs focuses on changes over several decades.
+(a) Exposure by education
+(b) Exposure by age and sex
+Figure 5: AI exposure by workers’ characteristics.
+(a) shows coefficients of regression of education categories on TEAI exposure.
+Controls include age, sex, occupation(4d), industry(3d), state and year fixed effects.
+(b) is a binscatter.
+The x-axis is the average age of workers in an industry-occupation-state observation in the 2022-18 ACS 5-year sample.
+Biscatter is computed considering education as a covariate.
+ACS individual weights are used.
 (a) TEAI index and employment growth
@@
 This figure plots the effect of AI score on employment and wage growth.
-Estimates are obtained from equation 4, with rolling regression coefficients and 95% confidence intervals of 4-year windows, starting in 2003-2007 and ending in 2019-2023.
-The point estimate refers to TEAI score, and the dependent variables are annual percentage growth rates of employment and wages.
+Estimates are from eq.
+3, with rolling regression coefficients and 95% confidence intervals of 4-year windows, starting in 2003-2007.
+The dependent variables are annual growth rates of employment and wages.
 Employment regression includes the log of the initial period of employment.
 Wage regression includes a log of initial period employment, log initial period wage and log initial period wage squared.
-All the regressions include occupation (SOC 4 digit) and sector (NAICS 3 digit) fixed effects.
-Robust standard errors clustered at NAICS level.
+Occupation and sector fixed effects included.
+Figure 7: Task exposure and task substitutability.
+Each dot represents an O*NET occupation.
+Division lines are constructed using median values of the distribution of both indexes.
+An interactive demo is available at https://terminatoreconomy.com/home
 5.3 Task Exposure or Substitutability?
@@
 As described in section 3 the TRAI index is specifically designed to assess task replacement.
-Figure 8: Task exposure and task substitutability
-This figure plots the TEAI index and the task replacement index TRAI .
-Each dot represents an O*NET occupation.
-Division lines are constructed using median values of the distribution of both indexes.
-An interactive visualisation and the full list of occupations and AI indexes are available at https://terminatoreconomy.com/home
-(a) Management, Business, Admin, Finance & ICT Science
-(b) Education, Social, Legal, Art & Health & Care
-(c) Other Services vs Sales
-(d) Farming & Fishing & Construction & Mining
-(e) Production & Repair &Transport
-Figure 9: Exposure and task substitutability by occupation groups.
-This figure plots the TEAI index and the replacement index TRAI .
-Each panel represents a pair of occupation groups.
-Each dot represents an occupation.
-Division lines are constructed using median values.
-What is the relationship between the TRAI and the TEAI index?
-On the one hand, the TRAI index can be considered as a sub-product of the TEAI index, as it is constructed by using the motivation summaries that LLMs produce in developing the TEAI .
-On the other hand, it differs because the task description that it analyzes is different (the TEAI uses ONET task description TRAI uses the summary of LMMs descriptions and motivations) and is designed to assess task substitutability, as the prompt asks LLM to specifically assess the possibility of replacing the human in task performance.
+What is the relationship between TRAI and TEAI index?
+On the one hand, the TRAI index can be considered a sub-product of the TEAI index, as it is constructed using the motivation summaries that LLMs produce in developing the TEAI.
+On the other hand, it differs because the task description that it analyzes is different (the TEAI uses ONET task description TRAI uses the summary of LLMs descriptions and motivations) and is designed to assess task substitutability, as the prompt asks LLM to specifically assess the possibility of replacing the human in task performance.
 Figure 7 displays the two indexes.
@@
 Also, while the figure shows different jobs, our replacement and exposure measures are constructed at task level.
-So, even if an occupation has a high level of replacement, it does not mean that it can be completely replaced by AI.
-Instead, it means that AI can perform a large number of tasks in that occupation, either completely or partially.
-The figure shows that the two measures are highly correlated, i.e.
-that high values of the TEAI index are generally associated with high values of the replacement index.
+Thus, even if an occupation has a high level of replacement, it does not mean that AI can completely replace it.
+Instead, it suggests that AI can perform a large number of tasks in that occupation, either completely or partially.
+The figure shows that the two measures are highly correlated, i.e., high values of the TEAI index are generally associated with high values of the replacement index.
 Grouping occupations improves the explanatory power of the analysis.
-In Figure 9 we isolate different occupation groups using the same taxonomy as in Figure 5(a).
 For example, most ICT and science occupations display a high level of exposure and task substitutability.
 This means that for these occupations, AI can execute several tasks with a high degree of confidence.
-Human activity will still be required, and it is likely that these professionals will concentrate on more critical and high value added tasks (such as supervision, analysis, developing solutions, etc.) leaving other tasks to AI.
-Analogously most teachers are characterized by a high level of exposure by a low level of substitutability: they are likely to work with AI while AI will replace few tasks.
-Similarly, several health and care occupations are characterized by low level of exposure and substitutability with few exceptions, such as “Medical record specialists” whose tasks are highly replaceable by technology.
+Human activity will still be required, and it is likely that these professionals will concentrate on more critical and high-value-added tasks (such as supervision, analysis, developing solutions, etc.), leaving other tasks to AI.
+Analogously, most teachers are characterized by a high level of exposure by a low level of substitutability: they are likely to work with AI, while AI will replace few tasks.
+Similarly, several health and care occupations are characterized by low levels of exposure and substitutability with few exceptions, such as “Medical record specialists” whose tasks are highly replaceable by technology.
 This provides some insights about the complementarity vs substitutability debate.
 Complementarity is often interpreted as AI and humans sharing the same activities.
-When looking at specific tasks, it emerges clearly that most of complementarity operates through substitution; in other words, some tasks will be executed by AI and machines while humans will concentrate on others.
-AI and humans will complement each other in the same occupation while the task allocation within occupation is likely to change.
-6 Conclusions
+When looking at specific tasks it emerges clearly that most of complementarity operates through substitution; in other words some tasks will be executed by AI and machines while humans will concentrate on others.
+AI and humans will complement each other in the same occupation, while task allocation within occupations is likely to change.
+6 Conclusions and Limitations
 We proposed a methodology for assessing AI exposure using LLMs, with a fine-grained approach that analyses exposure for 923 jobs and their 19281 tasks, using the O*NET US official taxonomy as a baseline.
@@
 We find that even in high-skill occupations, there is high variability in task substitution by AI, suggesting that AI and humans complement each other in the same occupation while allocating tasks within professions is likely to change.
-A human evaluation has been performed to assess the approach’s effectiveness, revealing that an average of 71-75% of the time evaluators agree with the score and explanations provided by our indexes, whilst only 2% of the time tend to disagree fully.
+A human evaluation has been performed to assess the approach’s effectiveness, revealing that an average of 71-75% of time evaluators agree with the score and explanations provided by our indexes, whilst only 2% of time tend to disagree fully.
 To ensure full reproducibility, our methodology relies only on top-performing open LLMs, and all occupation and task lists are accessible through an online demo.
 This would allow the community to track future assessments of improvements in new LLM models while integrating our findings to help individuals and organizations shape policies for AI integration into careers and the workforce.
-Finally, while O*NET offers a comprehensive and widely recognized taxonomy of the labour market, its static nature may not capture emerging tasks as they emerge in the real-labour market.
+Finally, while O*NET offers a comprehensive and widely recognized taxonomy of the labor market, its static nature may not capture emerging tasks as they emerge in the real-labor market.
 To address this limitation, our approach is designed to be flexible and adaptable to any user-defined taxonomy, enabling organizations to assess the impact of AI within their operational contexts accurately.
-Online Demo. We have made the results of our framework shown in this paper available at https://terminatoreconomy.com/home.
+Appendix A Appendix
+A.1 Evaluating Consensus Among LLMs
+To construct a single indicator, we took a conservative approach by assigning to each task the value of the rating with the highest frequency among the three models; if the three ratings differed, we chose the lowest.
+To assess the agreement between the three rates expressed by the LLMs, for each single task, we compute a consensus metric TASTLE2007531.
+$Cns(a_{i})=1+\sum_{k=1}^{m}p_{k}\log_{2}(1-\frac{\lvert LV_{k}-\mu_{LV}\rvert}{d_{LV}})$ (4)
+This process allows the three rates to be combined into a single measure of agreement.
+A higher consensus value indicates greater agreement between the observed rating values expressed for each task by the three different LLMs.
+The equation (4) shows the consensus calculation in which $LV_{k}$ represents the observed rating value, $p_{k}$ its relative frequency, $\mu_{LV}$ represents the weighted average of the $LV$ ratings using $p_{k}$ probabilities as weights, and $d_{L}V$ represents the scale size of the ratings adopted.
+The logarithmic function computes the impact of the normalised difference between each rating and the weighted average, moderated by the $d_{L}V$ dimension.
+The calculation uses a repeated summation for each k-th rate expressed for each individual task.
+Analogously, to estimate the similarity between the motivations provided by the LLMs, we compute the centroid of semantic cosine similarity rahutomo2012semantic, between them.
+The embedding vectors for the centroid computation is obtained using an open source Transformer model: as for the LLMs, we chose the Transformer model to be used in accordance with the Massive Text Embedding Benchmark (MTEB) Leaderboard by muennighoff-etal-2023-mteb.
+[^13] https://huggingface.co/spaces/mteb/leaderboard.
+Having English-language motivations, the choice fell on the UAE-Large-V1 https://huggingface.co/WhereIsAI/UAE-Large-V1, which represented an excellent compromise between effectiveness and efficiency, given its small size.
+In particular, since higher cosine similarity values reflect higher semantic similarities between the text of the LLM motivations, we expect a strong correlation between the consensus metric and cosine similarity.
+The heat map represented in Figure 8 shows that both the values of cosine similarity and the consensus metric are extremely high, with an average close to 0.9 in both cases.
+Figure 8: Heat map between cosine similarity of textual motivation of LLMs and consensus measure between scores. Data are aggregated at occupation level
+On the one hand, this suggests coherence between LLM-generated rates and the associated motivations, and on the other hand it adds robustness to our conservative approach in choosing between different models.
+Acknowledgements
+This paper is partially supported by the research activity of an EU H2020 Project PILLARS — Pathways to Inclusive Labor Markets - Grant agreement 101004703 - https://www.h2020-pillars.eu/
+Contribution Statement
+E.C., F.M., and M.M conceived the study; F.M., M.M., and A.S.
+developed the approach from a computational perspective E.C.
+performed the analyses from an economic perspective; F.M.
+managed the acquisition of funds.
+All authors wrote the initial manuscript draft, reviewed, and approved the final version of the manuscript.